# Orthogonal Subspace Decomposition for Generalizable AI-Generated Image Detection

### v2

@@ Abstract
AI-generated images (AIGIs), such as natural or face images, have become increasingly realistic and indistinguishable, making their detection a critical and pressing challenge. In this paper, we start from a new perspective to excavate the reason behind the failure generalization in AIGI detection, named the asymmetry phenomenon, where a naively trained detector tends to favor overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked, which is proved seriously limiting the expressivity and generalization. One potential remedy is incorporating the pre-trained knowledge within the vision foundation models (higher-ranked) to expand the feature space, alleviating the model’s overfitting to fake. To this end, we employ Singular Value Decomposition (SVD) to decompose the original feature space into two orthogonal subspaces. By freezing the principal components and adapting only the remained components, we preserve the pre-trained knowledge while learning forgery-related patterns. Compared to existing full-parameters and LoRA-based tuning methods, we explicitly ensure orthogonality enabling the higher rank of the whole feature space, effectively minimizing overfitting and enhancing generalization. Extensive experiments with our deep analysis on both deepfake and synthetic image detection benchmarks demonstrate superior generalization performance in detection.
@@ 1 Introduction
The rapid development of AI generative technologies has significantly lowered the barrier for creating highly realistic fake images. As deep generative models advance and mature (Goodfellow et al., 2020; Ho et al., 2020; Rombach et al., 2022b), the proliferation of AI-generated images (AIGIs [^1] In the context of this research, AIGI primarily refers to deepfakes (face-swapping) and synthetic images (e.g., nature or arts).) has drawn considerable attention, driven by their ability to produce high-quality content with relative ease. However, these advancements also introduce significant risks, if misused for malicious purposes such as deepfakes (mainly including face-swapping (Korshunov & Marcel, 2018) and face-reenactment (Thies et al., 2016)), which may violate personal privacy, spread misinformation, and erode trust in digital media. Consequently, there is an urgent need to develop a reliable and robust framework for detecting AIGIs.
@@
We have conducted extensive experiments on both deepfake detection and synthetic image detection benchmarks and surprisingly find that our approach achieves significant superiority over other SOTAs with very few training cost. Compared to existing full-parameters and LoRA-based tuning methods, we explicitly ensure orthogonality enabling the higher rank of the whole feature space, effectively minimizing overfitting to fake and enhancing generalization.
@@
  • Stricking generalization performance in detection: Effort achieves remarkably high generalization performance on previously unseen forgeries, which can be applied to detect a broad range of AIGIs, including but not limited to face forgeries (e.g., classical deepfakes) and synthetic natural images.
@@ 3 Methodology
Theoretical analysis for the asymmetric phenomena in AIGI Detection.
Here, we provide a theoretical explanation for the asymmetric phenomena in AIGI Detection.
Definition 3.1.
(Covariance Asymmetry Measure) Let $\Sigma_{\mathrm{real}}:=\mathbb{E}_{x\sim P_{\mathrm{real}}}[xx^{\top}]$ denote the covariance matrix of real data and $\Sigma_{\mathrm{fake}}:=\mathbb{E}_{z}[\mathbb{E}[x|z]\mathbb{E}[x|z]^{\top}]$ denote the covariance matrix of generator outputs, where $z$ denotes the latent variables. We define the asymmetry metric:
$\gamma:=\frac{\mathrm{rank}(\Sigma_{\mathrm{real}})}{\mathrm{rank}(\Sigma_{\mathrm{fake}})}-1$ (8)
Theorem 3.2.
Under the following assumptions: (i) The real data covariance satisfies $\mathrm{rank}(\Sigma_{\mathrm{real}})\geq r_{\mathrm{real}}$, (ii) Each generator $G_{j}:\mathbb{R}^{d}\to\mathbb{R}^{n}$ satisfies $\mathrm{rank}\left(\mathbb{E}[G_{j}(z)G_{j}(z)^{\top}]\right)\leq r_{\mathrm{fake}}$ for $j=1,\ldots,k$, then with probability exceeding $1-\delta$:
$\gamma\geq\frac{r_{\mathrm{real}}}{kr_{\mathrm{fake}}+C\sqrt{d\log(2k/\delta)}}-1,$ (9)
where $k$ is the number of generators, $d$ is the latent dimension, and $C$ is a universal constant.
This theory establishes the intrinsic asymmetry in AIGI detection through covariance spectral analysis, and uncovers that symmetric classifiers’ inevitable failure as $n/d\rightarrow\infty(\gamma\rightarrow\infty)$, providing mathematical grounding for explaining the observed asymmetric phenomena in AIGI detection. The proof can be seen in the appendix (Sec. A).
@@ 4.1 Deepfake Image Detection
Implementation Details. We utilize CLIP ViT-L/14 (Radford et al., 2021) as the default vision foundation model (VFM). We also investigate other VFMs in Tab. 5. We follow the pre-processing and training pipeline and use the codebases of DeepfakeBench (Yan et al., 2023b). Additionally, we sample 8 frames from each video for training and 32 frames for inference, following (Shiohara & Yamasaki, 2022). We use the fixed learning rate of 2e-4 for training our approach and employ the Adam (Kingma & Ba, 2014) for optimization. We set the batch size to 32 for both training and testing. We also employ several widely used data augmentations, such as Gaussian Blur and Image Compression, following other existing works (Yan et al., 2024a; Shiohara & Yamasaki, 2022; Cheng et al., 2024). For the evaluation metric, we report the widely-used video-level Area Under the Curve (AUC) to compare our approach with other works, following (Lin et al., 2024; Xu et al., 2023; Shiohara & Yamasaki, 2022). Similar to these works, we compute the average model’s output probabilities of each video to obtain the video-level AUC.
@@
To provide a comprehensive benchmark for comparison, we introduce 13 competitive detectors. All detectors are trained on FF++ (c23) and tested on other fake data. Results in Tab. 1 demonstrate two notable advantages of our approach. (1) generalizability: we see that our approach consistently and largely outperforms other models across basically all tested scenarios, validating that our method is generalizable for detecting unseen fake data, even for the latest face-swapping techniques such as BleFace (Shiohara et al., 2023). (2) efficiency: it is worth noting that our method only needs 0.19M parameters for training to achieve superior generalization. As we can see most latest SOTA detectors such as LSDA and ProDet all use about 100M parameters for training, while we are about 1,000$\times$ smaller.
@@ 4.3 Ablation Study and Analysis
Ablation studies in Tab. 4 demonstrate incremental gains from the proposed SVD method and loss constraints: the baseline (no modules) achieves an average AUC of 0.854, while adding SVD alone boosts performance to 0.909 (+6.5%), underscoring its efficacy in isolating forgery artifacts via orthogonal feature decomposition. Further integrating the singular value constraint ($\mathcal{L}_{\rm ksv}$) and orthogonal constraint ($\mathcal{L}_{\rm orth}$) refines performance, yielding 0.916 and 0.914 AUC, respectively, with their combined synergy (SVD + $\mathcal{L}_{\rm ksv}$ + $\mathcal{L}_{\rm orth}$) achieving peak performance (0.921 AUC). These results highlight the significant contribution of SVD as the primary driver of performance gains while illustrating the complementary benefits of $\mathcal{L}_{\rm ksv}$ and $\mathcal{L}_{\rm orth}$ in further refining the generalization performance.
@@
Figure 5: PCA quantifying feature space information retention, measured by the minimum number of components required to explain $\geq 90\%$ variance. Our SVD-based fine-tuning preserves 159/160 principal components (retaining 99.3% original variance), while LoRA and FFT exhibit notable information degradation, with 14.4% and 64.4% variance loss, respectively.
@@ 5 Conclusion
In this paper, we start our research from a new perspective to excavate the failure reason of the generalization in AIGI detection, namely the asymmetry phenomena, where a naively trained detector very quickly shortcuts to the seen fake patterns, collapsing the feature space into a low-ranked structure that limits expressivity and generalization. To address this, we propose integrating higher-ranked pre-trained knowledge from vision foundation models to expand the feature space. Simultaneously, we decompose the feature space into two orthogonal subspaces, for preserving pre-trained knowledge while learning forgery. Beyond LoRA and full-parameters tuning, we explicitly ensure the orthogonality maintaining the higher rank of the whole feature space for better generalization. We conduct extensive experiments with deep analysis on both deepfake and synthetic image detection benchmarks, demonstrating the superior advantages in generalization.
@@ 6 Impact Statements
Appendix A Proof of Theorem-3.2
Proof.
The covariance rank disparity emerges from distinct generation mechanisms. For real data covariance $\Sigma_{\mathrm{real}}$, the factor model assumption implies that $r_{\mathrm{real}}$ independent latent factors with non-degenerate interactions induce $\mathrm{rank}(\Sigma_{\mathrm{real}})\geq r_{\mathrm{real}}$ through singular value decomposition of the factor loading matrix.
For synthetic data, we decompose $\Sigma_{\mathrm{fake}}=\sum_{j=1}^{k}\lambda_{j}\Sigma_{j}+\Sigma_{\epsilon}$ where each component covariance $\Sigma_{j}=\mathbb{E}[G_{j}(z)G_{j}(z)^{\top}]$ satisfies $\mathrm{rank}(\Sigma_{j})\leq r_{\mathrm{fake}}$, and $\Sigma_{\epsilon}$ captures residual covariance. Through rank subadditivity, we obtain $\mathrm{rank}(\Sigma_{\mathrm{fake}})\leq\sum_{j=1}^{k}\mathrm{rank}(\Sigma_{j})+\mathrm{rank}(\Sigma_{\epsilon})\leq kr_{\mathrm{fake}}+d_{\epsilon}$.
Under the Lipschitz continuity condition $\|G_{j}(z)\|_{2}\leq L\|z\|_{2}$ for generators, we apply Gaussian chaos concentration bounds to control the residual term: $\mathbb{P}(\|\Sigma_{\epsilon}\|_{*}>C\sqrt{d\log(2k/\delta)})\leq\delta$. The union bound over $k$ generators consequently establishes the probabilistic rank bound, completing the proof. ∎
Appendix B Algorithm Procedure Illustration
Here, we provide an algorithm illustration of the proposed Effort approach in Alg. 1 for a comprehensive understanding.
Algorithm 1 Effort Approach Algorithm
0:  Updated weight matrix $W$
1:  $\triangleright$ Step 1: Singular Value Decomposition
2:  Decompose $W$ via SVD: $W=U\Sigma V^{\top}$
3:  Retain top $r$ singular values and vectors:
5:  Compute $W_{r}=U_{r}\Sigma_{r}V_{r}^{\top}$
6:  Keep $W_{r}$ fixed during training
7:  Compute residual component: $\Delta W=W-W_{r}$
8:  Decompose $\Delta W$ via SVD: $\Delta W=U_{n-r}\Sigma_{n-r}V_{n-r}^{\top}$
9:  Initialize $\Delta W$
10:  Define concatenated matrices:
11:   $\hat{U}=[U_{r},U_{n-r}]\in\mathbb{R}^{n\times n}$
12:   $\hat{V}=[V_{r},V_{n-r}]\in\mathbb{R}^{n\times n}$
13:  $\triangleright$ Step 2: Training Loop
14:  for each epoch do
15:    for each batch in $\mathcal{D}$ do
16:       $\triangleright$ Forward Pass
17:       Compute model output using $W=W_{r}+\Delta W$
18:       Compute classification loss $\mathcal{L}_{\rm cls}$
19:       $\triangleright$ Compute Constraints
20:       Compute orthogonality loss:
22:       Compute singular value constraint loss:
24:       $\triangleright$ Total Loss
26:       $\triangleright$ Backward Pass and Optimization
27:       Update $\Delta W$ using gradient descent to minimize $\mathcal{L}$
28:    end for
29:  end for
30:  $\triangleright$ Step 3: Output
31:  Update the weight matrix: $W\leftarrow W_{r}+\Delta W$
32:  return Updated weight matrix $W$
Appendix C Additional Results and Ablations
In this section, we provide additional experimental results and ablation studies to our proposed approach.
@@ Appendix C Additional Results and Ablations
C.1 Results on GenImage Benchmark
@@
C.2 Comparison with Existing Video Detectors
@@ C.3 Robustness Evaluation
Appendix D Additional cross-method evaluation
@@
Figure 10: t-SNE visualizations of the latent feature spaces between vanilla CNN (Wang et al., 2020b) and ours. We use the testing set of CycleGAN within UniversalFakeDetect Dataset (Wang et al., 2020b) for visualization.
D.1 Additional Ablation Studies
@@
Appendix E Additional Analysis and Visualizations
E.1 Additional t-SNE Visualizations
In our manuscript, we visualize the t-SNE of the seen fake ProGAN and unseen fake StyleGAN for the comparison of vanilla CNN (Res-50 (Wang et al., 2020b)) and ours (see Fig. 9). Now, we provide additional t-SNE visualizations, including t-SNE results between vanilla CNN and ours on CycleGAN, as shown in Fig. 10. As we can see from both two figures, our approach maximizes and preserves the pre-trained knowledge while fitting the forgery patterns during training, whereas the vanilla CNN overfits the seen fake method, learning forgery patterns only, thereby resulting in a highly low-ranked feature space (see Fig. 6 and Fig. 7 of the manuscript for details) and causing the overfitting to seen forgery patterns in the training set. Additionally, we see that the logit distribution of the vanilla CNN has a larger overlapping between fake and real, while ours is highly smaller, suggesting that our approach achieves a better generalization performance.
Appendix F Limitation and Future Work

### v4

@@ Abstract
AI-generated images (AIGIs), such as natural or face images, have become increasingly important yet challenging. In this paper, we start from a new perspective to excavate the reason behind the failure generalization in AIGI detection, named the asymmetry phenomenon, where a naively trained detector tends to favor overfitting to the limited and monotonous fake patterns, causing the feature space to become highly constrained and low-ranked, which is proved seriously limiting the expressivity and generalization. One potential remedy is incorporating the pre-trained knowledge within the vision foundation models (higher-ranked) to expand the feature space, alleviating the model’s overfitting to fake. To this end, we employ Singular Value Decomposition (SVD) to decompose the original feature space into two orthogonal subspaces. By freezing the principal components and adapting only the remained components, we preserve the pre-trained knowledge while learning fake patterns. Compared to existing full-parameters and LoRA-based tuning methods, we explicitly ensure orthogonality, enabling the higher rank of the whole feature space, effectively minimizing overfitting and enhancing generalization. We finally identify a crucial insight: our method implicitly learns a vital prior that fakes are actually derived from the real, indicating a hierarchical relationship rather than independence. Modeling this prior, we believe, is essential for achieving superior generalization. Our codes are publicly available at GitHub.
@@ 1 Introduction
The rapid development of AI generative technologies has significantly lowered the barrier for creating highly realistic fake images. As deep generative models advance and mature (Goodfellow et al., 2020; Ho et al., 2020; Rombach et al., 2022b; Yan et al., 2025), the proliferation of AI-generated images (AIGIs [^1] In the context of this research, AIGI primarily refers to deepfakes (face-swapping) and synthetic images (e.g., nature or arts).) has drawn considerable attention, driven by their ability to produce high-quality content with relative ease. However, these advancements also introduce significant risks, if misused for malicious purposes such as deepfakes (mainly including face-swapping (Korshunov & Marcel, 2018) and face-reenactment (Thies et al., 2016)), which may violate personal privacy, spread misinformation, and erode trust in digital media. Consequently, there is an urgent need to develop a reliable and robust framework for detecting AIGIs.
@@
We have conducted extensive experiments on both deepfake detection and synthetic image detection benchmarks and find that our approach achieves significant superiority over other SOTAs with very little training cost. Compared to existing full-parameters and LoRA-based tuning methods, we explicitly ensure orthogonality, enabling the higher rank of the whole feature space, effectively minimizing overfitting to fake and enhancing generalization.
Finally, we arrive at a key insight for generalizable AIGI detection: there exists an important prior of the detection task, where fake images are generated from real ones, establishing a hierarchical relationship rather than an independent or symmetric one. When aligning semantic information-such as distinguishing a fake dog from a real dog-this prior allows the model to focus discrimination within a smaller, semantically consistent subspace, e.g., only among dogs (see Fig. 4 for illustration). This focused discrimination simplifies the task and aligns with theoretical results from Rademacher complexity (Mohri & Rostamizadeh, 2008), which states that reducing model complexity leads to tighter generalization bounds. In contrast, naively trained detectors that treat real and fake data as independent fail to capture this structure, resulting in limited generalization performance. Therefore, modeling this hierarchical prior, we believe, is vitally crucial for AIGI detection.
@@
  • Key insight toward generalizable detection: We reveal that fake data is actually generated from real data, forming a hierarchical relationship rather than being independent. Our method effectively models this prior by maintaining the pre-trained semantic components while adapting to fake detection effectively, enabling the detector to make discrimination on the semantic-aligned subspaces, reducing model complexity and thus improving generalization.
Figure 4: t-SNE visualizations of the latent feature spaces between vanilla CNN (Wang et al., 2020b) and ours. Our method achieves orthogonal learning between the dimensions of semantic and forgery, allowing the model to capture fake patterns on the semantically-aligned subspace, simplifying the discrimination and thereby improving the generalization.
@@ 3 Methodology
Finally, we provide an algorithm illustration of the proposed approach in Alg. 1 for an overall understanding.
Algorithm 1 Effort Approach Algorithm
0:  Updated weight matrix $W$
1:  $\triangleright$ Step 1: Singular Value Decomposition
2:  Decompose $W$ via SVD: $W=U\Sigma V^{\top}$
3:  Retain top $r$ singular values and vectors:
5:  Compute $W_{r}=U_{r}\Sigma_{r}V_{r}^{\top}$
6:  Keep $W_{r}$ fixed during training
7:  Compute residual component: $\Delta W=W-W_{r}$
8:  Decompose $\Delta W$ via SVD: $\Delta W=U_{n-r}\Sigma_{n-r}V_{n-r}^{\top}$
9:  Initialize $\Delta W$
10:  Define concatenated matrices:
11:   $\hat{U}=[U_{r},U_{n-r}]\in\mathbb{R}^{n\times n}$
12:   $\hat{V}=[V_{r},V_{n-r}]\in\mathbb{R}^{n\times n}$
13:  $\triangleright$ Step 2: Training Loop
14:  for each epoch do
15:    for each batch in $\mathcal{D}$ do
16:       $\triangleright$ Forward Pass
17:       Compute model output using $W=W_{r}+\Delta W$
18:       Compute classification loss $\mathcal{L}_{\rm cls}$
19:       $\triangleright$ Compute Constraints
20:       Compute orthogonality loss:
22:       Compute singular value constraint loss:
24:       $\triangleright$ Total Loss
26:       $\triangleright$ Backward Pass and Optimization
27:       Update $\Delta W$ using gradient descent to minimize $\mathcal{L}$
28:    end for
29:  end for
30:  return Updated weight matrix $W\leftarrow W_{r}+\Delta W$
@@ 4.1 Deepfake Image Detection
Implementation Details. We utilize CLIP ViT-L/14 (Radford et al., 2021) as the default vision foundation model (VFM). We also investigate other VFMs in Tab. 5. We follow the pre-processing and training pipeline and use the codebases of DeepfakeBench (Yan et al., 2023b). Additionally, we sample 8 frames from each video for training and 32 frames for inference, following (Shiohara & Yamasaki, 2022). We use the fixed learning rate of 2e-4 for training our approach and employ the Adam (Kingma & Ba, 2014) for optimization. We set the batch size to 32 for both training and testing. We also employ several widely used data augmentations, such as Gaussian Blur and Image Compression, following other existing works (Yan et al., 2024a; Shiohara & Yamasaki, 2022; Cheng et al., 2024). For the evaluation metric, we report the widely-used video-level Area Under the Curve (AUC) to compare our approach with other works, following (Lin et al., 2024; Shiohara & Yamasaki, 2022). We compute the average model’s output probabilities of each video to obtain the video-level AUC.
@@
To provide a comprehensive benchmark for comparison, we introduce 13 competitive detectors, including several classical detection methods such as F3Net (Qian et al., 2020) (ECCV’20), SPSL (Liu et al., 2021a) (CVPR’20), SRM (Liu et al., 2021a) (CVPR’21), CORE (Ni et al., 2022) (CVPRW’22), RECCE (Cao et al., 2022) (CVPR’22), and SBI (Shiohara & Yamasaki, 2022) (CVPR’22), and also several latest SOTA methods (after 2023), such as UCF (Yan et al., 2023a) (ICCV’23), IID (Huang et al., 2023) (CVPR’23), TALL (Xu et al., 2023) (ICCV’23), LSDA (Yan et al., 2024a) (CVPR’24), ProDet (Cheng et al., 2024) (NeurIPS’24), and CFDA (Lin et al., 2024) (ECCV’24). All detectors are trained on FF++ (c23) and tested on other fake data. Results in Tab. 1 demonstrate two notable advantages of our approach. (1) generalizability: we see that our approach consistently and largely outperforms other models across basically all tested scenarios, validating that our method is generalizable for detecting unseen fake data, even for the latest face-swapping techniques such as BleFace (Shiohara et al., 2023). (2) efficiency: it is worth noting that our method only needs 0.19M parameters for training to achieve superior generalization. As we can see most latest SOTA detectors such as LSDA and ProDet all use about 100M parameters for training, while we are about 1,000$\times$ smaller.
@@ 4.3 Ablation Study and Analysis
Ablation studies in Tab. 4 demonstrate incremental gains from the proposed SVD method and loss constraints: the baseline (no modules) achieves an average AUC of 0.859, while adding SVD alone boosts performance to 0.925 (+6.6%), underscoring its efficacy in isolating forgery artifacts via orthogonal feature decomposition. Further integrating the singular value constraint ($\mathcal{L}_{\rm ksv}$) and orthogonal constraint ($\mathcal{L}_{\rm orth}$) refines performance, yielding 0.936 and 0.930 AUC, respectively, with their combined synergy (SVD + $\mathcal{L}_{\rm ksv}$ + $\mathcal{L}_{\rm orth}$) achieving peak performance (0.941 AUC). These results highlight the significant contribution of SVD as the primary driver of performance gains while illustrating the complementary benefits of $\mathcal{L}_{\rm ksv}$ and $\mathcal{L}_{\rm orth}$ in further refining the generalization performance.
@@
Figure 6: PCA quantifying feature space information retention, measured by the minimum number of components required to explain $\geq 90\%$ variance. Our SVD-based fine-tuning preserves 159/160 principal components (retaining 99.3% original variance), while LoRA and FFT exhibit notable degradation, with 14.4% and 64.4% variance loss, respectively. See Tab. 8 for more results.
@@
Table 8: Comparison of different fine-tuning methods. We compute both the effective rank and mean accuracy for evaluation. Our SVD retains most principal components (notably higher effective rank), thus achieving the best detection results.
Fine-tuning Methods Face Domain General Domain
Effective Rank Accuracy Effective Rank Accuracy
Baseline (No FT) 160 - 479 -
SVD (ours) 159 95.60 316 95.19
LoRA 137 89.40 304 93.03
FFT 57 85.70 238 86.22
@@
How to determine the value of rank in tuning SVD?
Our choice of using a lower rank (specifically, “n-r”=1) for fine-tuning DFD is primarily motivated by two critical factors: First, the nature of the real-fake classification task itself makes it relatively straightforward. Specifically, fake samples in existing training sets tend to exhibit a limited number of distinctive forgery patterns (FF++ contains only four forgery types), each with relatively simple and consistent characteristics. Due to this simplicity and limited diversity, a low-rank adaptation with a small rank (e.g., “n-r”=1, 4, or 16) is sufficient for the model to effectively learn these forgery patterns. As demonstrated by Table 5 in our paper, choosing ranks of 1, 4, or 16 yields very similar performance results. Given this observation, we prioritize efficiency and parameter economy, making rank 1 the optimal choice. Second, the inherent characteristics of binary classification further justify selecting a smaller rank. Binary classification tasks typically do not require the model to learn extensive and nuanced patterns, but rather to identify just enough distinctive features to separate the two classes, making the learned feature space inherently constrained. Thus, binary classification inherently simplifies the complexity of the learning problem, meaning that employing a higher rank would not provide significant additional benefit.
@@ 5 Conclusion
In this paper, we start our research from a new perspective to excavate the failure reason of the generalization in AIGI detection, namely the asymmetry phenomena, where a naively trained detector very quickly shortcuts to the seen fake patterns, collapsing the feature space into a low-ranked structure that limits expressivity and generalization. To address this, we propose integrating higher-ranked pre-trained knowledge from vision foundation models to expand the feature space. Simultaneously, we decompose the feature space into two orthogonal subspaces, for preserving pre-trained knowledge while learning forgery. Beyond LoRA and full-parameters tuning, we explicitly ensure the orthogonality, maintaining the higher rank of the whole feature space for better generalization. Furthermore, we reveal a very important prior for generalizable AIGI detection that fake data actually originates from real data in a hierarchical structure, not independently. Our method leverages this prior by preserving pre-trained semantic components while adapting to fake detection, enabling discrimination in semantic-aligned subspaces with reduced model complexity and improved generalization. Extensive experiments with deep analysis on both deepfake and synthetic image detection benchmarks have demonstrated the superior advantages of the proposed method in AIGI detection.
@@ 6 Impact Statements
7 Acknowledgment
This work was supported in part by the Natural Science Foundation of China (No. 62202014, 62332002, 62425101).
Appendix A Additional Results and Ablations
In this section, we provide additional experimental results and ablation studies of our proposed approach.
@@ Appendix A Additional Results and Ablations
A.1 Results on GenImage Benchmark
@@
A.2 Comparison with Existing Video Detectors
@@ A.3 Robustness Evaluation
Appendix B Additional cross-method evaluation
@@
B.1 Additional Ablation Studies
@@
Appendix C Additional Analysis and Visualizations
C.1 Additional t-SNE Visualizations
We further visualize the t-SNE of the seen fake ProGAN, unseen fake StyleGAN and useen CycleGAN for the comparison of vanilla CNN (Res-50 (Wang et al., 2020b)) and ours (see Fig. 9, Fig. 10). As we can see from both two figures, our approach maximizes and preserves the pre-trained knowledge while fitting the forgery patterns during training, whereas the vanilla CNN overfits the seen fake method, learning forgery patterns only, thereby resulting in a highly low-ranked feature space (see Fig. 6 and Fig. 7 of the manuscript for details) and causing the overfitting to seen forgery patterns in the training set. Additionally, we see that the logit distribution of the vanilla CNN has a larger overlapping between fake and real, while ours is highly smaller, suggesting that our approach achieves a better generalization performance.
C.2 Self-attention Map Visualizations
Here, we perform the self-attention maps visualization of the original CLIP-ViT model (Original), the fully fine-tuned CLIP-ViT model (FFT), the LoRA-trained CLIP-ViT model (LoRA), and our proposed orthogonal trained CLIP-ViT model (Ours) on the UniversalFakeDetect dataset (see Fig. 11, Fig. 12, Fig. 13 and Fig. 14). Specifically, for each block of the ViT, the self-attention map denotes the self-attention coefficient matrix calculated between the [CLS] token and the patch tokens. In the case of the LoRA component, the self-attention maps are generated from left to right using the original + LoRA weights, the original weights, and the LoRA weights, respectively. In the case of the Ours component, the self-attention maps are generated from left to right using the principal + residual weights, the principal weights, and the residual weights, respectively. Surprisingly, we observe that the semantic information is primarily concentrated in the earlier blocks, and our proposed approach establishes orthogonality between the semantic subspace and the learned forgery subspace at the level of the self-attention map. It further explains that our proposed approach can better preserve the pre-trained knowledge while learning fake patterns.
Figure 11: Self-attention map visualizations of UniversalFakeDetect Dataset (Wang et al., 2020b). We visualize the fake image of CycleGAN part and the corresponding real image for each block of the CLIP-ViT model (there are a total of 24 blocks, with IDs gradually increasing from top to bottom).
Figure 12: Self-attention map visualizations of UniversalFakeDetect Dataset (Wang et al., 2020b). We visualize the fake image of ProGAN part and the corresponding real image for each block of the CLIP-ViT model (there are a total of 24 blocks, with IDs gradually increasing from top to bottom).
Figure 13: Self-attention map visualizations of UniversalFakeDetect Dataset (Wang et al., 2020b). We visualize the fake image of DeepFake part and the corresponding real image for each block of the CLIP-ViT model (there are a total of 24 blocks, with IDs gradually increasing from top to bottom).
Figure 14: Self-attention map visualizations of UniversalFakeDetect Dataset (Wang et al., 2020b).We visualize the fake image of LDM part and the corresponding real image for each block of the CLIP-ViT model (there are a total of 24 blocks, with IDs gradually increasing from top to bottom).
Appendix D Limitation and Future Work